IMO training camp schedule 2002 Trinity College, Cambridge, 4-7 April

THURSDAY 4 APRIL 13.00-14.00: Arrival at Burrell’s Field 14.00-14.30: Ice-Breaker (ACM) 14.30-16.00: Introductory Problems (GCS) 16.00-16.30: Tea 16.30-18.15: Algebra (RA) 18.30-19.15: Dinner 19.30-21.15: Number Theory 1 (TG); Background Problems distributed (DM) late evening: Arrival of people who have been training for the Physics Olympiad.

FRIDAY 5 APRIL 08.15-08.45: Breakfast 09.00-10.45: Geometry 1 (DM) 10.45-11.15: Coffee 11.15-12.45: Inequalities (IBL) 13.00-14.00: Lunch 14.00-15.00: Trinity Talk and Prizes 15.00-16.00: Topology (VdS) 16.00-16.30: Tea 16.30-18.15: Combinatorics 1 (CCF) 18.30-19.15: Dinner 19.30-21.30: Relays (RA)

SATURDAY 6 APRIL 08.15-08.45: Breakfast 09.00-10.45: Functional Equations (ASa) / Geometry 2 (CB) / Comb. 2 (GCS) 10.45-11.15: Coffee 11.15-12.45: Geometric Inequalities (ACM) / Comb. 2 (GCS) / Number Theory 2 (TG) 13.00-14.00: Lunch 16.00-16.30: Tea 16.30-18.15: Geometry 2 (CB) / Number Theory 2 (TG) / Problem Solving (GCS) 18.30-19.15: Dinner 19.30-21.00: Reception 21.00-22.00: Staff Meeting

SUNDAY 7 APRIL 07.30-08.00: Breakfast 08.15-12.45: First Selection Test 10:00-12:00: BMOC Executive Committee Meeting 13.00-14.00: Lunch 14.00-15.30: Background Problems discussed (DM) 15.30-16.00: Tea and departure from Burrell’s Field 16.00-19.00: FST Marking